Find the slope of the ling PQ.
P is (13,2) and Q is (-5,26)

Write the equation of the line through the points P and Q in the form of y=mx+b.

Answer by Osran_Shri(18)

Point P ( x=13, y=2) Point Q ( x1= -5, y1 = 26)
In y = mx + b equation Slope of Line PQ is :
m = Delta ( Y ) / Delta (x)

m = (y1 - y)/(x1-x) = (26-2) / (-5-13) = -(24/18)= -1.3333 ( Note : negative slope )
Therefore Slope m = -1.333
Now Constant b = Y intercept i.e., Point where line PQ crosses Y axis at x=0
y = mx +b
b = y - mx = y1-mx1
b = 2 - (-1.333)(13) = 2 + 17.329 = 19.329 = 19.33( rounded off to 2 digits after decimal point)
Therefore : Foa a line PQ : Slope m = -1.333 and b = 19.33
Now Equation can be written in standard format as below :
y = mx + b
y = -1.333x + 19.33
Thats all...
B.Shridhar